### v2

@@ Appendix B Dutch Auctions (Section 3) Appendix
So far we have proven that $\mu(V,\xi)=\left\{\left(W_{\mathsf{L}},\tilde{\vec{\xi}}_{\mathsf{L}}\right),\left(W_{\mathsf{H}},\tilde{\vec{\xi}}_{\mathsf{H}}\right)\right\}.$ To complete the proof, we have to prove that $\tilde{\vec{\xi}}_{\mathsf{L}}=\vec{\xi}_{\mathsf{L}},\tilde{\vec{\xi}}_{\mathsf{H}}=\vec{\xi}_{\mathsf{H}}$. Towards contradiction, suppose $\tilde{\vec{\xi}}_{\mathsf{L}}\neq\vec{\xi}_{\mathsf{L}}$ or $\tilde{\vec{\xi}}_{\mathsf{H}}\neq\vec{\xi}_{\mathsf{H}}$. If $\tilde{\vec{\xi}}_{\mathsf{H}}\neq\vec{\xi}_{\mathsf{H}}$, then, by Lemma A.5, Condition (ii), there exists $i$ such that $\overline{{V}}_{i}=\overline{{V}}_{\xi},\left(\overline{{V}}_{i},i\right)\rhd\left(\overline{{V}}_{\xi},\xi\right)$. We can let $R=\left\{i\right\}$ and then apply Lemma B.4 to contradict the hypothesis that the auction is weakly shill-proof. If $\tilde{\vec{\xi}}_{\mathsf{L}}\neq\vec{\xi}_{\mathsf{L}}$, then, as argued in the proof of Theorem 3.4, the menu presented to $\tilde{\vec{\xi}}_{\mathsf{L}}$ must not have the auction end immediately, no matter what partition $\tilde{\vec{\xi}}_{\mathsf{L}}$ selects. Thus, our previous argument for the case where $\tilde{\vec{\xi}}_{\mathsf{H}}\neq\vec{\xi}_{\mathsf{H}}$ applies, and we can conclude that $\mu(V,\xi)=\mu^{D}(V,\xi)$. ∎
Proof of Theorem 3.6
@@
The statement follows as a corollary of Lemma B.5. Consider any optimal reserve $\rho^{*}$, $\underline{M}$ atoms below the optimal reserve, and $\overline{M}$ atoms above the optimal reserve. We construct a sparse (and regular) distribution $\tilde{F}$ with optimal reserve $\rho^{*}$, $\underline{M}$ atoms below the optimal reserve, and $\overline{M}$ atoms above the optimal reserve. To begin, let $\delta$ such that $\underline{M}\delta\leq\rho^{*}$ and $(\underline{M}+1)\delta>\rho^{*}$. Then for all $k\leq\underline{M}$, let $v^{k+1}-v^{k}=\delta$ and $\tilde{f}(v^{k})=e^{-\lambda(k-2)\delta}-e^{-\lambda(k-1)\delta}.$ Note that
$\tilde{\varphi}^{k}=v^{k}-(v^{k+1}-v^{k})\frac{1-\tilde{F}(v^{k})}{\tilde{f}(v^{k})}=(k-1)\delta-\delta\frac{e^{-\lambda(k-1)\delta}}{e^{-\lambda(k-2)\delta}-e^{-\lambda(k-1)\delta}},$
@@
In order for $\rho^{*}$ to be an optimal reserve of $\tilde{F}$, it must be the case that for $k^{*}$ such that $\tilde{\varphi}^{k^{*}}\geq 0$ and $\tilde{\varphi}^{k^{*}-1}<0$, it is also the case that $\rho^{*}\in((k^{*}-1)\delta,k^{*}\delta]$. Such a $k^{*}$ must be equal to $\lceil\check{k}\rceil$, where $\check{k}\delta-\frac{\delta}{e^{\lambda\delta}-1}=0.$ Thus, $\underline{M}-1=k^{*}=\lceil\frac{1}{e^{\lambda\delta}-1}\rceil$.
@@
Finally, to finish constructing $\tilde{F}$, we simply select atoms $v^{\underline{M}+2},\ldots,v^{\underline{M}+\overline{M}}$ and respective probability weights to satisfy
$\tilde{\varphi}^{k}\text{ is non-decreasing and }\sum_{k=\underline{M}+2}^{\underline{M}+\overline{M}}\tilde{f}(v^{k})=e^{-\lambda(\underline{M}+1)\delta}.$
This system of constraints has at most $\overline{M}$ constraints and $2(\overline{M}-1)$ free variables, so the system can be satisfied. Thus, we have constructed a regular and sparse $\tilde{F}$ that has the required values of $\rho^{*}$,$\underline{M}$, and $\overline{M}$. Then, we can apply Lemma B.5 to conclude the proof.
Appendix C Weakly Shill-Proof and Strategy-Proof Auctions (Section 4) Appendix
@@
For such a distribution $F$, let $\overline{\Delta}=\max_{k}\left\{v^{k}-v^{k-1}\right\}$. As convention, let $F^{-1}$ be the left pseudo-inverse: $F^{-1}(x)=\max\left\{v^{k}:x\geq F(v^{k})\right\}$.
@@
$v^{Y}\geq F^{-1}\left(F\left(\rho^{*}\right)+\max_{1\leq n<N}\left\{\left(\max\left\{1-\frac{\rho^{*}}{\rho^{*}+2\overline{\Delta}}\left(\frac{f(\rho^{*})}{1-F(\rho^{*})}\right)^{n},0\right\}\right)^{1/n}\right\}\right),$ (6)
the ascending, screening auction with screening level $v^{Y}$ is a weakly shill-proof, ex-post incentive compatible, and optimal auction.
@@
The Ascending, Screening Auction is Orderly and Optimal. We first prove that the auction is well-defined, orderly, and optimal. The transfer and allocation function are orderly and optimal, so we only have to show that the menu rule can induce this outcome function. Let us examine the English auction phase first. The auction ends if and only if $\overline{{V}}<v^{M}$. When that occurs, the auction has determined $v_{i}$ for all $i$ given $v_{i}\geq\rho^{*}$. Thus, the outcome is fully determined. In the second-price auction phase, each value weakly greater than $v^{Y}$ is determined precisely (and there are at least two players with values weakly greater than $v^{Y}$) and so the outcome rule is determined.
The Ascending, Screening Auction is Ex-Post Incentive Compatible. Observe that the definition of ex-post incentive compatiblility (Definition 4.1) is a function solely of the direct mechanism $(\tilde{x},\tilde{t})$ and not of the menu rule $\mu$. Both the English auction phase and the second-price auction use the same transfer function $\tilde{t}^{2}$. The entire auction has the optimal allocation rule $\tilde{x}^{*}$ and so the ascending, screening auction is ex-post incentive compatible.
The Ascending, Screening Auction is Weakly Shill-Proof. By assumption that the screen level is at least $\rho^{*}$, any potential shill bidder will be asked to play at least once in the English auction before being able to play in the second-price auction. If the optimal shill bid is 00 in the first round of the English auction, then the auction is weakly shill-proof because once a bidder reports 00, she “drops out” and does not take another action.
Observe that given the form of the transfer rule, the maximum amount that a bidder $i$ with value $v_{i}$ would have to pay is $v_{i}$. Thus, the maximum possible gain in revenue from a shill bidder deviating is at most the difference between the first and second moment of $v$. Next, note that MHR distributions are regular. Regularity implies that if a shill bidder reports a non-zero value in the English auction stage and the auction concludes before reaching the second-price stage, the expected gain must be weakly less than 00. So, when considering the expected gain of misreporting, we can think of the expected gain from manipulating outcomes in the English auction component as at most 00 and can focus on manipulating outcomes in the second-price stage. Therefore, the total gains from misreporting as a shill bidder must be bounded above by the probability that a shill bidder is able to manipulate the outcome of the second-price auction multiplied by the expected difference between the first and second moments of the value distribution conditional on reaching the second-price auction stage.
@@
Suppose bidder $i$ is a shill bidder and it is the first time she is taking an action. Then, under the rules of the auction, she has not indicated that her value is greater than $\rho^{*}$ yet. For any real bidder $j\neq i$, there are two cases: either bidder $j$ has indicated her value is weakly greater than $\rho^{*}$ $\left(\mathds{P}\left[v_{j}<v^{Y}\right]=F(v^{Y})-F(\rho^{*})\right)$ or she has not yet taken an action $\left(\mathds{P}\left[v_{j}<v^{Y}\right]=F(v^{Y})\right)$. So, if $K\leq N$ real bidders have not dropped out yet (i.e., indicated that their value is less than $\rho^{*}$), then the probability that the auction would continue to the second-price auction is at most $1-\left(F\left(v^{Y}\right)-F\left(\rho^{*}\right)\right)^{K}$. Therefore, the maximum expected gain for a shill bidder from misreporting in her first action of the English auction phase when $K$ bidders have not dropped is at most
@@
$\mathds{P}\left[v_{j}\leq v^{m}\right]\geq\frac{\displaystyle\sum_{\left\{k:\underline{{V}}_{j}\leq v^{k}<v^{m}\right\}}f(v^{k})}{1-F(\underline{{V}}_{j})}\geq\frac{f(\underline{{V}}_{j})}{1-F(\underline{{V}}_{j})}\geq\frac{f(\rho^{*})}{1-F(\rho^{*})}.$
@@
$\rho^{*}\cdot\left(\frac{f(\rho^{*})}{1-F(\rho^{*})}\right)^{K}.$ (8)
We conclude the proof by showing that $v^{Y}$ satisfying Equation 6 implies that the expected revenue loss from misreporting as a shill is weakly larger than the expected gain. Beginning with Equation 6, we can see that for all $K<N$,
$v^{Y}$ $\geq F^{-1}\left(F\left(\rho^{*}\right)+\max_{1\leq n<N}\left\{\left(\max\left\{1-\frac{\rho^{*}}{\rho^{*}+2\overline{\Delta}}\left(\frac{f(\rho^{*})}{1-F(\rho^{*})}\right)^{n},0\right\}\right)^{1/n}\right\}\right)$
$\geq F^{-1}\left(F\left(\rho^{*}\right)+\left(\max\left\{1-\frac{\rho^{*}}{\rho^{*}+2\overline{\Delta}}\left(\frac{f(\rho^{*})}{1-F(\rho^{*})}\right)^{K},0\right\}\right)^{1/K}\right)$
@@
The left-hand side of the preceding equation corresponds to Equation 8, the lower bound on the expected loss from misreporting as a shill bidder, and the right-hand side corresponds to Equation 7, the upper bound on the expected gain from misreporting and thus we have shown that it is equilibrium not to shill when $v^{Y}$ is sufficiently high. ∎
Proof of Theorem 4.5.
Let $F_{m}$ be the discrete approximation of the exponential distribution with rate $\lambda=1$ and atoms at $\left\{0,2,\ldots,2m\right\}$. Then, using the argument from the proof of Theorem 3.6, $F_{m}$ is regular and MHR. For all $m>2$, an optimal reserve is $\rho^{*}=4$. Observe that $Y^{*}=2$ satisfies Equation 6 because
$\frac{\rho^{*}}{\rho^{*}+2\overline{\Delta}}\left(\frac{f(\rho^{*})}{1-F(\rho^{*})}\right)^{n}=\frac{1}{2}(e^{2}-1)^{n}>1\text{ for all }n\geq 1.$
We apply Lemma C.3 to conclude that the ascending, screening auction with screen level $v^{Y^{*}}$ is weakly shill-proof, ex-post incentive compatible, and optimal for all $F_{m}$. Then,
$Q^{AS,Y^{*}}(F_{m})=N\left(Y^{*}-\min\left\{k:\rho^{*}\leq v^{k}\right\}+2\right)=2N\text{ and}$
$Q^{E}(F_{m})=N\left(m-\min\left\{k:\rho^{*}\leq v^{k}\right\}+1\right)=(m-1)N.$
Thus, $Q^{AS}(F_{m},Y^{*})/Q^{E}(F_{m})\to 0$ as $m\to\infty$, concluding the proof. ∎
Appendix D Single-Action Auctions (Section 5) Appendix

### v3

@@ Appendix C Weakly Shill-Proof Auctions (Section 4) Appendix
So far we have proven that $\mu(\Theta,\xi)=\left\{\left(W_{\mathsf{L}},\tilde{\vec{\xi}}_{\mathsf{L}}\right),\left(W_{\mathsf{H}},\tilde{\vec{\xi}}_{\mathsf{H}}\right)\right\}.$ To complete the proof, we have to prove that $\tilde{\vec{\xi}}_{\mathsf{L}}=\vec{\xi}_{\mathsf{L}}$ and $\tilde{\vec{\xi}}_{\mathsf{H}}=\vec{\xi}_{\mathsf{H}}$. Towards contradiction, suppose $\tilde{\vec{\xi}}_{\mathsf{L}}\neq\vec{\xi}_{\mathsf{L}}$ or $\tilde{\vec{\xi}}_{\mathsf{H}}\neq\vec{\xi}_{\mathsf{H}}$. If $\tilde{\vec{\xi}}_{\mathsf{H}}\neq\vec{\xi}_{\mathsf{H}}$, then, by Lemma A.4, Condition (ii), there exists $i$ such that $\overline{{\Theta}}_{i}=\overline{{\Theta}}_{\xi},\left(\overline{{\Theta}}_{i},i\right)\rhd\left(\overline{{\Theta}}_{\xi},\xi\right)$. We can let $R=\left\{i\right\}$ and then apply Lemma C.6 to contradict the hypothesis that the auction is weakly shill-proof. If $\tilde{\vec{\xi}}_{\mathsf{L}}\neq\vec{\xi}_{\mathsf{L}}$, then, as argued in the proof of Theorem 3.6, the menu presented to $\tilde{\vec{\xi}}_{\mathsf{L}}$ must not have the auction end immediately, no matter what partition $\tilde{\vec{\xi}}_{\mathsf{L}}$ selects. Thus, our previous argument for the case where $\tilde{\vec{\xi}}_{\mathsf{H}}\neq\vec{\xi}_{\mathsf{H}}$ applies, and we can conclude that $\mu(\Theta,\xi)=\mu^{D}(\Theta,\xi)$. ∎
Proof of Proposition 4.6
@@
The statement follows as a corollary of Lemma C.7. Consider any optimal reserve $\theta^{\rho^{*}}$, $\underline{M}$ atoms below the optimal reserve, and $\overline{M}$ atoms above the optimal reserve. We construct a sparse (and regular) distribution $\tilde{F}$ with optimal reserve $\theta^{\rho^{*}}$, $\underline{M}$ atoms below the optimal reserve, and $\overline{M}$ atoms above the optimal reserve. To begin, let $\delta$ such that $\underline{M}\delta\leq\theta^{\rho^{*}}$ and $(\underline{M}+1)\delta>\theta^{\rho^{*}}$. Then for all $k\leq\underline{M}$, let $\theta^{k+1}-\theta^{k}=\delta$ and $\tilde{f}(\theta^{k})=e^{-\lambda(k-2)\delta}-e^{-\lambda(k-1)\delta}.$ Note that
$\tilde{\varphi}^{k}=\theta^{k}-(\theta^{k+1}-\theta^{k})\frac{1-\tilde{F}(\theta^{k})}{\tilde{f}(\theta^{k})}=(k-1)\delta-\delta\frac{e^{-\lambda(k-1)\delta}}{e^{-\lambda(k-2)\delta}-e^{-\lambda(k-1)\delta}},$
@@
In order for $\theta^{\rho^{*}}$ to be an optimal reserve of $\tilde{F}$, it must be the case that for $k^{*}$ such that $\tilde{\varphi}^{k^{*}}\geq 0$ and $\tilde{\varphi}^{k^{*}-1}<0$, it is also the case that $\theta^{\rho^{*}}\in((k^{*}-1)\delta,k^{*}\delta]$. Such a $k^{*}$ must be equal to $\lceil\check{k}\rceil$, where $\check{k}\delta-\frac{\delta}{e^{\lambda\delta}-1}=0.$ Thus, $\underline{M}-1=k^{*}=\lceil\frac{1}{e^{\lambda\delta}-1}\rceil$.
@@
Finally, to finish constructing $\tilde{F}$, we simply select atoms $\theta^{\underline{M}+2},\ldots,\theta^{\underline{M}+\overline{M}}$ and respective probability weights to satisfy
$\tilde{\varphi}^{k}\text{ is non-decreasing and }\sum_{k=\underline{M}+2}^{\underline{M}+\overline{M}}\tilde{f}(\theta^{k})=e^{-\lambda(\underline{M}+1)\delta}.$
This system of constraints has at most $\overline{M}$ constraints and $2(\overline{M}-1)$ free variables, so the system can be satisfied. Thus, we have constructed a regular and sparse $\tilde{F}$ that has the required values of $\theta^{\rho^{*}}$$\underline{M}$, and $\overline{M}$. Then, we can apply Lemma C.7 to conclude the proof.
C.3 Strategy-Proof Auctions (Section 4.2.2) Appendix
@@
For such a distribution $F$, let $\overline{\Delta}=\max_{k}\left\{\theta^{k}-\theta^{k-1}\right\}$. As convention, let $F^{-1}$ be the left pseudo-inverse: $F^{-1}(x)=\max\left\{\theta^{k}:x\geq F(\theta^{k})\right\}$.
@@
$\theta^{Y}\geq F^{-1}\left(F\left(\theta^{\rho^{*}}\right)+\max_{1\leq n<N}\left\{\left(\max\left\{1-\frac{\theta^{\rho^{*}}}{\theta^{\rho^{*}}+2\overline{\Delta}}\left(\frac{f(\theta^{\rho^{*}})}{1-F(\theta^{\rho^{*}})}\right)^{n},0\right\}\right)^{1/n}\right\}\right),$ (12)
the ascending, screening auction with screening level $\theta^{Y}$ is a weakly shill-proof, ex-post incentive compatible, and optimal auction.
@@
The Ascending, Screening Auction is Orderly and Optimal. We first prove that the auction is well-defined, orderly, and optimal. The transfer and allocation function are orderly and optimal, so we only have to show that the menu rule can induce this outcome function. Let us examine the English auction phase first. The auction ends if and only if $\overline{{\Theta}}<\theta^{M}$. When that occurs, the auction has determined $\theta_{i}$ for all $i$ given $\theta_{i}\geq\theta^{\rho^{*}}$. Thus, the outcome is fully determined. In the second-price auction phase, each value weakly greater than $\theta^{Y}$ is determined precisely (and there are at least two players with values weakly greater than $\theta^{Y}$) and so the outcome rule is determined.
@@
The Ascending, Screening Auction is Weakly Shill-Proof. By assumption that the screen level is at least $\theta^{\rho^{*}}$, any potential shill bidder will be asked to play at least once in the English auction before being able to play in the second-price auction. If the optimal shill bid is 0 in the first round of the English auction, then the auction is weakly shill-proof because once a bidder reports 0, she “drops out” and does not take another action.
Observe that given the form of the transfer rule, the maximum amount that a bidder $i$ with value $\theta_{i}$ would have to pay is $\theta_{i}$. Thus, the maximum possible gain in revenue from a shill bidder deviating is at most the difference between the first and second moment of $F$. Next, note that MHR distributions are regular. Regularity implies that if a shill bidder reports a non-zero value in the English auction stage and the auction concludes before reaching the second-price stage, the expected gain must be weakly less than 0. So, when considering the expected gain of misreporting, we can think of the expected gain from manipulating outcomes in the English auction component as at most 0 and can focus on manipulating outcomes in the second-price stage. Therefore, the total gains from misreporting as a shill bidder must be bounded above by the probability that a shill bidder is able to manipulate the outcome of the second-price auction multiplied by the expected difference between the first and second moments of the value distribution conditional on reaching the second-price auction stage.
[…]
Suppose bidder $i\in S$ and it is the first time she is taking an action. Then, under the rules of the auction, she has not indicated that her value is greater than $\theta^{\rho^{*}}$ yet. For any real bidder $j\neq i$, there are two cases: either bidder $j$ has indicated her value is weakly greater than $\theta^{\rho^{*}}$ $\left(\mathds{P}\left[\theta_{j}<\theta^{Y}\right]=F(\theta^{Y})-F(\theta^{\rho^{*}})\right)$ or she has not yet taken an action $\left(\mathds{P}\left[\theta_{j}<\theta^{Y}\right]=F(\theta^{Y})\right)$. So, if $K\leq N$ real bidders have not dropped out yet (i.e., indicated that their value is less than $\theta^{\rho^{*}}$), then the probability that the auction would continue to the second-price auction is at most $1-\left(F\left(\theta^{Y}\right)-F\left(\theta^{\rho^{*}}\right)\right)^{K}$. Therefore, the maximum expected gain for a shill bidder from misreporting in her first action of the English auction phase when $K$ bidders have not dropped is at most
@@
$\mathds{P}\left[\theta_{j}\leq\theta^{m}\right]\geq\frac{\displaystyle\sum_{\left\{k:\underline{{\Theta}}_{j}\leq\theta^{k}<\theta^{m}\right\}}f(\theta^{k})}{1-F(\underline{{\Theta}}_{j})}\geq\frac{f(\underline{{\Theta}}_{j})}{1-F(\underline{{\Theta}}_{j})}\geq\frac{f(\theta^{\rho^{*}})}{1-F(\theta^{\rho^{*}})}.$
@@
$\theta^{\rho^{*}}\cdot\left(\frac{f(\theta^{\rho^{*}})}{1-F(\theta^{\rho^{*}})}\right)^{K}.$ (14)
We conclude the proof by showing that $\theta^{Y}$ satisfying Equation 12 implies that the expected revenue loss from misreporting as a shill is weakly larger than the expected gain. Beginning with Equation 12, we can see that for all $K<N$,
$\theta^{Y}$ $\geq F^{-1}\left(F\left(\theta^{\rho^{*}}\right)+\max_{1\leq n<N}\left\{\left(\max\left\{1-\frac{\theta^{\rho^{*}}}{\theta^{\rho^{*}}+2\overline{\Delta}}\left(\frac{f(\theta^{\rho^{*}})}{1-F(\theta^{\rho^{*}})}\right)^{n},0\right\}\right)^{1/n}\right\}\right)$
$\geq F^{-1}\left(F\left(\theta^{\rho^{*}}\right)+\left(\max\left\{1-\frac{\theta^{\rho^{*}}}{\theta^{\rho^{*}}+2\overline{\Delta}}\left(\frac{f(\theta^{\rho^{*}})}{1-F(\theta^{\rho^{*}})}\right)^{K},0\right\}\right)^{1/K}\right)$
@@
The left-hand side of the preceding equation corresponds to Equation 14, the lower bound on the expected loss from misreporting as a shill bidder, and the right-hand side corresponds to Equation 13, the upper bound on the expected gain from misreporting and thus we have shown that it is equilibrium not to shill when $\theta^{Y}$ is sufficiently high. ∎
Proof of Proposition 4.12.
Let $F_{m}$ be the discrete approximation of the exponential distribution with rate $\lambda=1$ and atoms at $\left\{0,2,\ldots,2m\right\}$. Then, using the argument from the proof of Proposition 4.6, $F_{m}$ is regular and MHR. For all $m>2$, an optimal reserve is $\theta^{\rho^{*}}=4$. Observe that $Y^{*}=3$ satisfies Equation 12 because
$\frac{\theta^{\rho^{*}}}{\theta^{\rho^{*}}+2\overline{\Delta}}\left(\frac{f(\theta^{\rho^{*}})}{1-F(\theta^{\rho^{*}})}\right)^{n}=\frac{1}{2}(e^{2}-1)^{n}>1\text{ for all }n\geq 1.$
We apply Lemma C.10 to conclude that the ascending, screening auction with screen level $\theta^{Y^{*}}$ is weakly shill-proof, ex-post incentive compatible, and optimal for all $F_{m}$. Then, $Q^{AS,Y^{*}}(F_{m})=2$ and $Q^{E}(F_{m})=m-2$. Thus, $Q^{AS}(F_{m},Y^{*})/Q^{E}(F_{m})\to 0$ as $m\to\infty$, concluding the proof. ∎
Appendix D Affiliation and Interdependence (Section 5) Appendix
Lemma D.1 (Karlin and Rinott (1980)).
For any non-decreasing function $g$, if $F^{\prime}\succeq_{\mathrm{Aff}}F$, then $\mathds{E}_{\theta\sim F^{\prime}}\left[g(\theta)\right]\geq\mathds{E}_{\theta\sim F}\left[g(\theta)\right]$.